# Channel Dependence, Limited Lookback Windows, and the Simplicity of Datasets: How Biased is Time Series Forecasting?

### v2

@@ Appendix I Cell Cycle ODE Discussion
Table 21. Hyperparameter settings for the ETTm1 dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:0.0002,e_layers:7,d_ff:1.02e+03,d_model:1.02e+03,dropout:0.648,fc_dropout:0.175,patch_size:8,stride:4,seq_len:512
PatchTST CD-96 lr:1.60e-06,e_layers:8,d_ff:256,d_model:1.02e+03,dropout:0.307,fc_dropout:0.623,patch_size:8,stride:4,seq_len:512
@@
PatchTST CI-192 lr:7.02e-07,e_layers:2,d_ff:1.02e+03,d_model:512,dropout:0.292,fc_dropout:0.673,patch_size:8,stride:4,seq_len:336
PatchTST CD-192 lr:1.82e-06,e_layers:4,d_ff:1.02e+03,d_model:1.02e+03,dropout:0.471,fc_dropout:0.405,patch_size:16,stride:8,seq_len:512
@@
PatchTST CI-336 lr’: 0.0002, ’e_layers’: 7, ’d_ff’: 1024, ’d_model’: 1024, ’dropout’: 0.6482, ’fc_dropout’: 0.1748, ’patch_size’: 8, ’stride’: 4, ’seq_len’: 512
PatchTST CD-336 lr’: 1.602e-06, ’e_layers’: 8, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.307, ’fc_dropout’: 0.6229, ’patch_size’: 8, ’stride’: 4, ’seq_len’: 512
@@
PatchTST CI-720 lr:2.27e-06,e_layers:4,d_ff:1.02e+03,d_model:128,dropout:0.198,fc_dropout:0.356,patch_size:16,stride:8,seq_len:512
PatchTST CD-720 lr:0.0032,e_layers:3,d_ff:1.02e+03,d_model:128,dropout:0.318,fc_dropout:0.402,patch_size:16,stride:8,seq_len:512
@@
Table 22. Hyperparameter settings for the ETTm2 dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
@@
PatchTST CI-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.014, ’fc_dropout’: 0.539, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
PatchTST CD-336 lr:0.0004,num_blocks:8,hidden_size:1.02e+03,dropout:0.159,activation:relu,seq_len:512
@@
PatchTST CD-720 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
TSMixer CI-720 lr:0.0002,num_blocks:8,hidden_size:1.02e+03,dropout:0.205,activation:relu,seq_len:512
@@
Table 23. Hyperparameter settings for the Hopfield dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:0.0003,e_layers:9,d_ff:256,d_model:128,dropout:0.228,fc_dropout:0.072,patch_size:8,stride:8,seq_len:336
@@
PatchTST CI-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
PatchTST CD-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
@@
PatchTST CI-336 lr’: 0.0003, ’e_layers’: 9, ’d_ff’: 256, ’d_model’: 128, ’dropout’: 0.2277, ’fc_dropout’: 0.0716, ’patch_size’: 8, ’stride’: 8, ’seq_len’: 336
PatchTST CD-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.5389, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
@@
Table 24. Hyperparameter settings for the Lorenz dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
@@
PatchTST CI-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
PatchTST CD-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
@@
PatchTST CI-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.539, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
PatchTST CD-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.5389, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
@@
PatchTST CI-720 lr:0.0001,e_layers:2,d_ff:1.02e+03,d_model:1.02e+03,dropout:0.463,fc_dropout:0.708,patch_size:16,stride:4,seq_len:192
PatchTST CD-720 lr:0.0018,e_layers:5,d_ff:1.02e+03,d_model:128,dropout:0.097,fc_dropout:0.623,patch_size:16,stride:8,seq_len:512
@@
Table 25. Hyperparameter settings for the LorenzCoupled dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96
@@
TSMixer CI-96 lr:0.0067,num_blocks:9,hidden_size:1.02e+03,dropout:0.318,activation:relu,seq_len:512
@@
TSMixer CI-192 lr:0.0067,num_blocks:9,hidden_size:1.02e+03,dropout:0.318,activation:relu,seq_len:512
@@
TimeMixer CI-192 lr 0.00008, d_ff 512, d_model 512, e_layers 4, seq_len 192
@@
PatchTST CI-336 lr:0.0067,num_blocks:9,hidden_size:1.02e+03,dropout:0.318,activation:relu,seq_len:512
@@
CrossFormer CD-336 lr’: 0.00017626081318364395, ’e_layers’: 7, ’d_ff’: 128, ’d_model’: 512, ’dropout’: 0.3870070203430017, ’seg_len’: 10, ’baseline’: 0, ’cross_factor’: 16, ’seq_len’: 192, ’affine’: 1
@@
Table 26. Hyperparameter settings for the electricity dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:0.0003,e_layers:9,d_ff:256,d_model:128,dropout:0.228,fc_dropout:0.072,patch_size:8,stride:8,seq_len:336
PatchTST CD-96 lr:0.0001,e_layers:1,d_ff:1.02e+03,d_model:128,dropout:0.098,fc_dropout:0.465,patch_size:8,stride:8,seq_len:336
TSMixer CI-96 lr:0.0002,num_blocks:5,hidden_size:1.02e+03,dropout:0.205,activation:relu,seq_len:512
@@
PatchTST CD-192 lr:0.0002,e_layers:1,d_ff:512,d_model:1.02e+03,dropout:0.010,fc_dropout:0.333,patch_size:16,stride:8,seq_len:192
TSMixer CI-192 lr:0.0002,num_blocks:5,hidden_size:1.02e+03,dropout:0.247,activation:relu,seq_len:512
@@
TimeMixer CI-192 lr 0.00008, d_ff 512, d_model 512, e_layers 4, seq_len 192
@@
TSMixer CI-336 lr:0.0002,num_blocks:5,hidden_size:1.02e+03,dropout:0.265,activation:relu,seq_len:512
@@
PatchTST CD-720 lr:6.07e-06,e_layers:3,d_ff:256,d_model:1.02e+03,dropout:0.006,fc_dropout:0.642,patch_size:16,stride:8,seq_len:192
TSMixer CI-720 lr:0.0003,num_blocks:2,hidden_size:1.02e+03,dropout:0.108,activation:relu,seq_len:512
@@
Table 27. Hyperparameter settings for the weather dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96
@@
CrossFormer CI-96 lr’: 0.000515701560401399, ’e_layers’: 7, ’d_ff’: 128, ’d_model’: 256, ’dropout’: 0.25730612136617825, ’seg_len’: 11, ’baseline’: 0, ’cross_factor’: 10, ’seq_len’: 96, ’affine’: 1
CrossFormer CD-96 lr’: 5.2122436504604685e-05, ’e_layers’: 3, ’d_ff’: 256, ’d_model’: 512, ’dropout’: 0.12303593071645767, ’seg_len’: 8, ’baseline’: 0, ’cross_factor’: 7, ’seq_len’: 336, ’affine’: 0
@@
TSMixer CI-192 lr:0.0067,num_blocks:9,hidden_size:1.02e+03,dropout:0.318,activation:relu,seq_len:512
@@
TSMixer CI-336 lr:0.0067,num_blocks:9,hidden_size:1.02e+03,dropout:0.318,activation:relu,seq_len:512
@@
Table 28. Hyperparameter settings for the Traffic dataset. The best variant for each model is shown.

### v3

@@ Appendix 0.J Ablation Study on RevIN
Hyperparameter settings for the ETTm2 dataset across different forecasting horizons.
Model Setting Best Hyperparameters
[…]
Model Setting Best Hyperparameters
[…]
PatchTST CD-96 lr: 1.60e-06, e_layers: 8, d_ff: 256, d_model: 1024, dropout: 0.307, fc_dropout: 0.623, patch_size: 8, stride: 4, seq_len: 512
@@
PatchTST CI-192 lr: 7.02e-07, e_layers: 2, d_ff: 1024, d_model: 512, dropout: 0.292, fc_dropout: 0.673, patch_size: 8, stride: 4, seq_len: 336
PatchTST CD-192 lr: 1.82e-06, e_layers: 4, d_ff: 1024, d_model: 1024, dropout: 0.471, fc_dropout: 0.405, patch_size: 16, stride: 8, seq_len: 512
@@
PatchTST CI-336 lr: 0.0002, e_layers: 7, d_ff: 1024, d_model: 1024, dropout: 0.6482, fc_dropout: 0.1748, patch_size: 8, stride: 4, seq_len: 512
PatchTST CD-336 lr: 1.60e-06, e_layers: 8, d_ff: 256, d_model: 1024, dropout: 0.307, fc_dropout: 0.6229, patch_size: 8, stride: 4, seq_len: 512
@@
PatchTST CI-720 lr: 2.27e-06, e_layers: 4, d_ff: 1024, d_model: 128, dropout: 0.198, fc_dropout: 0.356, patch_size: 16, stride: 8, seq_len: 512
PatchTST CD-720 lr: 0.0032, e_layers: 3, d_ff: 1024, d_model: 128, dropout: 0.318, fc_dropout: 0.402, patch_size: 16, stride: 8, seq_len: 512
@@
Hyperparameter settings for the Hopfield dataset across different forecasting horizons.
Model Setting Best Hyperparameters
[…]
Model Setting Best Hyperparameters
[…]
@@
PatchTST CI-336 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
PatchTST CD-336 lr: 0.0004, num_blocks: 8, hidden_size: 1024, dropout: 0.159, activation: relu, seq_len: 512
@@
PatchTST CD-720 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
TSMixer CI-720 lr: 0.0002, num_blocks: 8, hidden_size: 1024, dropout: 0.205, activation: relu, seq_len: 512
@@
Hyperparameter settings for the Lorenz dataset across different forecasting horizons.
Model Setting Best Hyperparameters
[…]
Model Setting Best Hyperparameters
[…]
@@
PatchTST CI-192 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
PatchTST CD-192 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
@@
PatchTST CI-336 lr: 0.0003, e_layers: 9, d_ff: 256, d_model: 128, dropout: 0.2277, fc_dropout: 0.0716, patch_size: 8, stride: 8, seq_len: 336
PatchTST CD-336 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.0137, fc_dropout: 0.5389, patch_size: 16, stride: 8, seq_len: 192
@@
Hyperparameter settings for the LorenzCoupled dataset.
Model Setting Best Hyperparameters
[…]
Model Setting Best Hyperparameters
[…]
@@
PatchTST CI-192 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
PatchTST CD-192 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
@@
PatchTST CI-336 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.0137, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
PatchTST CD-336 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.0137, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
@@
PatchTST CI-720 lr: 0.0001, e_layers: 2, d_ff: 1024, d_model: 1024, dropout: 0.463, fc_dropout: 0.708, patch_size: 16, stride: 4, seq_len: 192
PatchTST CD-720 lr: 0.0018, e_layers: 5, d_ff: 1024, d_model: 128, dropout: 0.097, fc_dropout: 0.623, patch_size: 16, stride: 8, seq_len: 512
@@
Hyperparameter settings for the electricity dataset.
Model Setting Best Hyperparameters
[…]
Model Setting Best Hyperparameters
[…]
@@
TSMixer CI-96 lr: 0.0067, num_blocks: 9, hidden_size: 1024, dropout: 0.318, activation: relu, seq_len: 512
@@
TSMixer CI-192 lr: 0.0067, num_blocks: 9, hidden_size: 1024, dropout: 0.318, activation: relu, seq_len: 512
@@
TimeMixer CI-192 lr: 0.00008, d_ff: 512, d_model: 512, e_layers: 4, seq_len: 192
@@
PatchTST CI-336 lr: 0.0067, num_blocks: 9, hidden_size: 1024, dropout: 0.318, activation: relu, seq_len: 512
@@
CrossFormer CD-336 lr: 0.000176, e_layers: 7, d_ff: 128, d_model: 512, dropout: 0.387, seg_len: 10, baseline: 0, cross_factor: 16, seq_len: 192, affine: 1
@@
Hyperparameter settings for the weather dataset.
Model Setting Best Hyperparameters
[…]
Model Setting Best Hyperparameters
[…]
PatchTST CD-96 lr: 0.0001, e_layers: 1, d_ff: 1024, d_model: 128, dropout: 0.098, fc_dropout: 0.465, patch_size: 8, stride: 8, seq_len: 336
TSMixer CI-96 lr: 0.0002, num_blocks: 5, hidden_size: 1024, dropout: 0.205, activation: relu, seq_len: 512
@@
PatchTST CD-192 lr: 0.0002, e_layers: 1, d_ff: 512, d_model: 1024, dropout: 0.010, fc_dropout: 0.333, patch_size: 16, stride: 8, seq_len: 192
TSMixer CI-192 lr: 0.0002, num_blocks: 5, hidden_size: 1024, dropout: 0.247, activation: relu, seq_len: 512
@@
TimeMixer CI-192 lr: 0.00008, d_ff: 512, d_model: 512, e_layers: 4, seq_len: 192
@@
TSMixer CI-336 lr: 0.0002, num_blocks: 5, hidden_size: 1024, dropout: 0.265, activation: relu, seq_len: 512
@@
PatchTST CD-720 lr: 6.07e-06, e_layers: 3, d_ff: 256, d_model: 1024, dropout: 0.006, fc_dropout: 0.642, patch_size: 16, stride: 8, seq_len: 192
TSMixer CI-720 lr: 0.0003, num_blocks: 2, hidden_size: 1024, dropout: 0.108, activation: relu, seq_len: 512
@@
Hyperparameter settings for the Traffic dataset.
Model Setting Best Hyperparameters
[…]
Model Setting Best Hyperparameters
[…]
@@
CrossFormer CI-96 lr: 0.000516, e_layers: 7, d_ff: 128, d_model: 256, dropout: 0.257, seg_len: 11, baseline: 0, cross_factor: 10, seq_len: 96, affine: 1
CrossFormer CD-96 lr: 5.21e-05, e_layers: 3, d_ff: 256, d_model: 512, dropout: 0.123, seg_len: 8, baseline: 0, cross_factor: 7, seq_len: 336, affine: 0
@@
TSMixer CI-192 lr: 0.0067, num_blocks: 9, hidden_size: 1024, dropout: 0.318, activation: relu, seq_len: 512
@@
TSMixer CI-336 lr: 0.0067, num_blocks: 9, hidden_size: 1024, dropout: 0.318, activation: relu, seq_len: 512